# Emergent Weyl-like points in periodically modulated systems and Floquet engineering

### v3

@@ VII Summary
In our study, we introduce a novel 3D topological phase with Weyl-semimetal-like characteristics, which is induced by a periodic potential. This phase is realized by stacking 2D Chern insulators, leading to Weyl-like points in the parameter space $(k_{x},k_{y},z)$, in contrast to the conventional Weyl points in momentum space $(k_{x},k_{y},k_{z})$.
@@
The Weyl-semimetal-like phase we propose shares key features with traditional Weyl semimetals, such as linear dispersion near the Weyl-like points, nontrivial bulk topology, and the emergence of Fermi arc that connect these points. Importantly, these topological behaviors manifest in real space, rather than in momentum space.
@@
Moreover, we investigate the effects of interlayer coupling by calculating the local density of states, the layer Hall conductance, and the total 3D Hall conductivity. Our results show that the Weyl-semimetal-like phase remains robust under weak and moderate interlayer couplings.
Additionally, our investigation explores the effects of high-frequency periodic driving, where we calculate Floquet Weyl nodes, Floquet Chern numbers, Floquet energy spectra, and Floquet Fermi arcs. We demonstrate that the Weyl-semimetal-like nodes can be tuned via high-frequency laser driving, providing a means to manipulate these topological features. This work opens new possibilities for exploring and controlling topological phases through external stimuli.
@@ Acknowledgements.
We acknowledge helpful discussions with Ching Hua Lee. F.Q. acknowledges the support from the Jiangsu Specially-Appointed Professor Program in Jiangsu Province and the Doctoral Research Start-Up Fund of Jiangsu University of Science and Technology. R.C. acknowledges the support from the National Natural Science Foundation of China (Grant No. 12304195) and the Chutian Scholars Program in Hubei Province.

### v4

@@ IX Summary
In our study, we introduce a 3D topological phase with Weyl-semimetal-like characteristics, which is induced by a periodic potential. This phase is realized by stacking 2D Chern insulators, leading to Weyl-like points in the parameter space $(k_{x},k_{y},z)$, in contrast with the conventional Weyl points in momentum space $(k_{x},k_{y},k_{z})$.
@@
The Weyl-semimetal-like phase we propose shares key features with traditional Weyl semimetals, such as linear dispersion near the Weyl-like points, nontrivial bulk topology, the emergence of Fermi arc that connect these points, and the Berry monopoles. Importantly, these topological behaviors manifest in real space rather than in momentum space.
@@
We further analyze the role of interlayer coupling by computing the local density of states, the layer Hall conductance, and the total 3D Hall conductivity and show that the Weyl-semimetal-like phase remains robust under weak and moderate coupling. The effects of disorder are also investigated: Beyond a critical disorder strength, the Weyl-like points destabilize, driving the system into a trivial phase.
Additionally, we explore Floquet engineering under high-frequency periodic driving, calculating Floquet Weyl nodes, Floquet Chern numbers, Floquet energy bands, and Floquet Fermi arcs. Our results demonstrate that the Weyl-like nodes can be effectively tuned by laser fields, offering a route to manipulate topological properties dynamically.
Finally, we reveal the emergence of both type-I and II Weyl-like behaviors in a tilted Weyl-like model. In this work, we open possibilities for exploring and controlling topological phases through external stimuli.
@@ Acknowledgements.
We acknowledge helpful discussions with Ching Hua Lee. F.Q. acknowledges support from the Jiangsu Specially-Appointed Professor Program in Jiangsu Province and the Doctoral Research Start-Up Fund of Jiangsu University of Science and Technology. R.C. acknowledges support from the National Natural Science Foundation of China (Grant No. 12304195), the Chutian Scholars Program in Hubei Province, the Hubei Provincial Natural Science Foundation (Grant No. 2025AFA081), and the Original Seed Program of Hubei University.
Data Availability
The data are available from the authors upon reasonable request.